OpenStudy (anonymous):

A boat requires four hours to travel a given distance upstream against a five mile per hour current, but only requires three hours to travel downstream with that same current back to the dock. Find the speed of the boat in still water. PLEASE HELP!! Thanks

OpenStudy (anonymous):

Hi jason it's Diana, Melissa just asked this one too

OpenStudy (anonymous):

Seems like we are all in here.....lol

OpenStudy (anonymous):

I'm stuck on number 8

OpenStudy (anonymous):

let v be the speed of boat when boat travels against,v-5=d/4 v+5=d/3 solving gives,v=7d/24

OpenStudy (anonymous):

If it helps at all, this is what I came up with for this one.............. Boat speed = v Current = 5 mph Upstream against the current = v - 5 Downstream with the current = v + 5 Time traveling upstream = 4 hrs Time traveling downstream = 3 hrs 4(v - 5) = 3(v + 5) 4v - 20 = 3v + 15 - 3v - 3v v - 20 = 15 + 20 +20 v = 35 The speed of the boat in still water is 35 mph

OpenStudy (anonymous):

AWESOME! Thanks foe everyone help!

OpenStudy (anonymous):

np, did you get number 8 yet?

OpenStudy (anonymous):

8? Nickels and quarters question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well I am Confused? Are you in my math class?

OpenStudy (anonymous):

Class LOL

OpenStudy (anonymous):

Well I am Confused? Are you in my math class?

OpenStudy (anonymous):

yes